OpenStudy (decarr432):

if f(x)=5x-4 and g(x)=x^2, what is (g^of)(2)

OpenStudy (anonymous):

first find \(f(2)\)

OpenStudy (anonymous):

then whatever you get, find \(g(\text{that})\)

OpenStudy (decarr432):

How do I find f(2)?

OpenStudy (anonymous):

not to be mean, but if you cannot find \(f(2)\) then you certainly cannot find \((g\circ f)(2)\)

OpenStudy (decarr432):

I am the worst at math ;-;

OpenStudy (anonymous):

you evaluate a function at a number by plugging in the number where you see the variable

OpenStudy (anonymous):

for example, if \(f(x)=2x+4\) then \(f(5)=2\times 5+4=14\)

OpenStudy (decarr432):

Okay so simple plug (2) in for all my x's

OpenStudy (anonymous):

right you have \(f(x)=x^2\) right?

OpenStudy (decarr432):

g(x)=x^2

OpenStudy (anonymous):

ooh ok and \[f(x)=5x-4\]

OpenStudy (decarr432):

yeah

OpenStudy (anonymous):

so if \[f(x)=5x-4\] what is \(f(2)\)?

OpenStudy (decarr432):

it would be 5(2)-4=6 right and x^2 would be the same as 2*2

OpenStudy (anonymous):

wait slow down

OpenStudy (anonymous):

you are right, \[f(2)=5\times 2-4=6\]

OpenStudy (anonymous):

but now you do not want to compute \(g(2)=2^2=4\) instead you want to compute \(g(6)=6^2\)

OpenStudy (decarr432):

okay because 6 is now my new x so I plug it instead of 2. right?

OpenStudy (anonymous):

yes

OpenStudy (decarr432):

6^2=36?

OpenStudy (anonymous):

in one line it looks like this \[(g\circ f)(2)=g(f(2))=g(6)=36\]

OpenStudy (anonymous):

yes 36 is your answer

OpenStudy (decarr432):

Okay thank you sorry for so many questions, I did horrible in match in middle school and its coming to bite me in the butt in highschool lol

OpenStudy (anonymous):

that wasn't so hard, was it?

OpenStudy (anonymous):

no problem we can do another if you iike

OpenStudy (decarr432):

No it wasn't thank you so much for the help

OpenStudy (decarr432):

Ill post the next one thanks